You would like to know the average hourly wage in your area. You ask 32 people, and find a sample mean of $10.75. The population standard deviation is $3.69. (Assume that the distribution is normal.)




Round your answers to 2 decimal places




a.) Find the Point Estimate for the mean hourly wage



b.) Find the 95% margin of error for the mean

1
Expert's answer
2021-12-07T10:30:12-0500

a) Point estimate = 10.75

b) Z for 95% confidence interval = Z = 1.96

Margin of error "= Z \times \frac{sd }{ \sqrt{32}} = 1.96 \times \frac{3.69}{\sqrt{32}} = 1.278"